OpenStudy (anonymous):

Check please!! vv 3x-[5-4(2x-7)] my answer was x-7

OpenStudy (anonymous):

also I have...\[\sqrt{16x^2}\times \sqrt{3x^3}\] my answer was...\[4x^2\sqrt{3x}\]

OpenStudy (anonymous):

then the first answer is x+7 ?

OpenStudy (anonymous):

first I did parentheses 5-4 = 1 then keeping in parentheses distribute -1 to 2x -7 then 3x-2x =x and (-1)(-7) = 7....right??

OpenStudy (anonymous):

3x - [5 - 4 (2x - 7)] The most inside parentheses first so distribute the 4 3x - [5 - 8x - 28]

OpenStudy (anonymous):

if you are distributing -4 times -7 wouldn't it be +28 ?

OpenStudy (anonymous):

Yup.

OpenStudy (anonymous):

then would you just add/subtract like terms?

OpenStudy (anonymous):

3x - [5 - 8x + 28] combine like terms inside parentheses 3x - [ - 8x + 33] distribute negative 3x + 8x - 33

OpenStudy (anonymous):

okay thank you!! Can you also check my work for my second problem way up there?

OpenStudy (anonymous):

Is that sqrt (16x^2) * sqrt (3x^2)? I can't really read it.

OpenStudy (anonymous):

If it's sqrt (16x^2) * sqrt (3x^3) then it's good.

OpenStudy (anonymous):

yes it's sqrt 3x^3 thank you both!

OpenStudy (anonymous):

:)